# Arithmetic Properties of Several Generalized-Constant Sequences, with Implications for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$

### v4

@@ Abstract.
Neither the Euler-Mascheroni constant, $\gamma=0.577215\ldots$, nor the Euler-Gompertz constant, $\delta=0.596347\ldots$, is currently known to be irrational. However, it has been proved that at least one of them is transcendental. The two constants are related through a well-known equation of Hardy, equivalent to $\gamma+\delta/e=\textrm{Ein}\left(1\right)$, which recently has been generalized to $\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e=\eta^{\left(n\right)},\>n\geq 0$ for sequences of constants $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ (given respectively by raw, conditional, and partial moments of the Gumbel(0,1) probability distribution). Investigating the $\gamma^{\left(n\right)}$ through recurrence relations (where $\gamma^{\left(0\right)}=1$ and $\gamma^{\left(1\right)}=\gamma$), we find that at least one of the pair $\{\gamma,\gamma^{\left(2\right)}\}$ and – conditional on a realistic conjecture verified for $2\leq n\leq 26$ – at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often (with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$). We then show, via a theorem of Shidlovskii, that the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$, implying that at least one of each pair, $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\}$ and $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\}$, and at least two of the triple $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\}$, are transcendental for all $n\geq 1$. Further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ reveals that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often with positive asymptotic density. Finally, we provide parallel results for the sequences $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$.
@@ 1.3. Outline of Results
In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with the $\gamma^{\left(n\right)}$. These recurrences reveal that at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and – conditional on a realistic conjecture computationally verified for $2\leq n\leq 26$ – at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental, implying that the $\gamma^{\left(n\right)}$ are transcendental infinitely often (with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$). Next, in Section 3, we apply a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$. This implies that at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$. Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$. Finally, in Section 4 we consider two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provide results parallel to those for the original system.
@@ 2.2.1. Algebraic Structure of the $P_{n}$
Assume (for purposes of contradiction) that the polynomials $\left\{P_{2},P_{3},\ldots,P_{n}\right\}$ are algebraically dependent; that is, there exists a non-zero polynomial $Q_{n}\in\mathbb{Q}\left[t_{2},t_{3},\ldots,t_{n}\right]$ such that $Q_{n}\left(P_{2},P_{3},\ldots,P_{n}\right)=0$. Writing $Q_{n}$ as a polynomial in $t_{n}$ gives
@@ 2.3. Transcendence Results
The transformation of the original polynomials of (16) into the reduced form of (17) lays the foundation for part (ii) of Theorem 1 below. However, one additional condition on the $\widetilde{P}_{n,\theta}$ is required. Although we know there exists $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ that can be expressed as a non-trivial polynomial equation in $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ of the form given by (19), it remains possible that $A_{d}^{\left(n,\theta\right)}=0$ for all $d\in\left\{1,2,\ldots,D_{n,\theta}\right\}$ at the actual realizations of the elements $\gamma^{\left(k\right)}\in\mathcal{G}_{n}^{\left(2n\right)}$ (to be denoted by $\boldsymbol{\gamma^{*}}\Longleftrightarrow\gamma^{\left(k\right)}=\textrm{E}_{X}\left[X_{\textrm{G}}^{k}\right]\>\forall\>k\in\mathbb{Z}_{\geq 1}$). Lemma 5 below, supported by Conjecture 1, ensures this does not happen.
@@
The conjecture below, which we have verified computationally for all $n\in\left\{2,3,\ldots,26\right\}$, asserts that condition (21) of Lemma 5 is true. [^8] For each $n$, we executed a Python program (on an Apple M4 Max MacBook computer with 48 GB memory) to provide rigorous certification of the sufficient condition in (21). The program first evaluated $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ symbolically for each $m\in\left\{2,3,\ldots,n\right\}$, and then confirmed $\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ using interval arithmetic. The maximum value of $n$ tested (i.e., $26$) was simply the largest $n$ for which the program completed all checks within approximately 24 hours. Clearly, greater values of $n$ could be tested, but the program’s runtime increases approximately exponentially in $n$. (For example, $n=30$ would require a runtime of about 1 week.) In addition to computational verification, the conjecture is realistic on theoretical grounds because it holds for any given $n$ unless $\boldsymbol{\gamma^{*}}$ is forced to lie on one of finitely many special hypersurfaces (i.e., the vanishing loci of the minors $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$).
@@
(iv) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ (i.e., $\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $\Omega\left(\ln\left(N+2\right)/N\right)$ as $N\rightarrow\infty$.
@@
(iv) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, we can identify a lower bound on the frequency with which transcendental values $\gamma^{\left(n\right)}$ can arise by constructing the sparsest feasible subsequence of transcendental elements $\left\{\gamma^{\left(\ell_{i}\right)}\right\}_{i\geq 1}$ satisfying parts (i) and (ii) as follows:
$\bullet\>$ Let $\ell_{1}=1$ and $\ell_{2}=4$.
$\bullet\>$ Given a set of previous element indices $\left\{\ell_{1},\ell_{2},\ldots,\ell_{i-1}\right\}$, select the new element index, $\ell_{i}$, to be as large as possible while still including $\ell_{i-1}+1$ in the interval $\left[\left\lceil\ell_{i}/2\right\rceil,\ell_{i}\right]$ (i.e., $\ell_{i}=2\ell_{i-1}+2$).
This procedure generates the sequence $\left\{\ell_{i}\right\}_{i\geq 1}=\left\{1,4,10,\ldots,3\cdot 2^{i-1}-2,\ldots\right\}$, implying that out of the first $3\cdot 2^{i-1}-2$ positive integers there are at least $i$ element indices – or equivalently, out of the first $N$ positive integers there are at least $\left\lfloor\log_{2}\left(\left(N+2\right)/3\right)+1\right\rfloor=\left\lfloor\ln\left(\left(N+2\right)/3\right)/\ln\left(2\right)+1\right\rfloor$ transcendental elements. As $N\rightarrow\infty$, an asymptotic lower bound on the density of the set of transcendental $\gamma^{\left(n\right)}$ is thus
$\Omega\left(\dfrac{1}{N}\left[\dfrac{\ln\left(N+2\right)-\ln\left(3\right)}{\ln\left(2\right)}+1\right]\right)=\Omega\left(\dfrac{\ln\left(N+2\right)}{N}\right).$
From the $\textrm{Gumbel}\left(0,1\right)$ moment-generating function, $\textrm{M}_{X}\left(t\right)=\Gamma\left(1-t\right),\>t<1$, it is easy to see that
@@
so that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere. In particular, if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]). [^9] It may be of interest to some readers to note that the asymptotic lower bound on the density of the set of posited transcendental $\Gamma^{\left(n\right)}\left(1\right)$ is smaller (in magnitude) than the asymptotic lower bound on the density of the set of irrational odd-$n$ $\zeta\left(n\right)$ given by Fischler, Sprang, and Zudilin [12]: $\Omega\left(\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{N/2}\right)=\Omega\left(N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1}\right),$ for arbitrarily small $\varepsilon>0$.
@@ 3. Algebraic Independence of the $\eta^{\left(n\right)}$
From (16), one can see that for any $m\in\left\{2,3,\ldots,n\right\}$, $P_{m}$ is linear in the highest-index constant, $\gamma^{\left(2m\right)}$, with a non-zero integer coefficient $\alpha_{m,0}$. This implies that each $\gamma^{\left(2m\right)}$ can be expressed as a rational polynomial in strictly lower-index constants; that is, $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}\right)$. We now let
@@
For $m=2$, the relation $P_{2}=0$ expresses $\gamma^{\left(4\right)}$ as a rational polynomial in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)}\in\mathcal{B}_{n}$, implying $\gamma^{\left(4\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$. Now consider $m\geq 3$, and assume that $\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$. Since (a) $P_{m}=0$ expresses $\gamma^{\left(2m\right)}$ as a rational polynomial in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}$, (b) every odd-indexed element of the set $\left\{\gamma^{\left(3\right)},\gamma^{\left(5\right)},\ldots,\gamma^{\left(2m-1\right)}\right\}$ lies in $\mathcal{B}_{n}$, and (c) every even-indexed element of $\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\right\}$ lies in $\mathbb{Q}\left(\mathcal{B}_{n}\right)$, it follows that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ as well. This confirms that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ for all $m\in\left\{2,3,\ldots,\lfloor n/2\rfloor\right\}$, and thus that $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}=\mathbb{Q}\left(\mathcal{G}_{n}\right)=\mathbb{Q}\left(\mathcal{B}_{n}\right)$.
@@
Consequently, as $N\rightarrow\infty$, we have the following asymptotic lower bound on the density of the set of transcendental $\delta^{\left(n\right)}/e$:

### v5

@@ Abstract.
Neither the Euler-Mascheroni constant, $\gamma=0.577215\ldots$, nor the Euler-Gompertz constant, $\delta=0.596347\ldots$, is currently known to be irrational. However, it has been proved that at least one of them is transcendental. The two constants are related through a well-known equation of Hardy, equivalent to $\gamma+\delta/e=\textrm{Ein}\left(1\right)$, which recently has been generalized to $\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e=\eta^{\left(n\right)},\>n\geq 0$ for sequences of constants $\gamma^{\left(n\right)}$, $\delta^{\left(n\right)}$, and $\eta^{\left(n\right)}$ (given respectively by raw, conditional, and partial moments of the Gumbel(0,1) probability distribution). Investigating the $\gamma^{\left(n\right)}$ through recurrence relations (where $\gamma^{\left(0\right)}=1$ and $\gamma^{\left(1\right)}=\gamma$), we find that at least one of the pair $\{\gamma,\gamma^{\left(2\right)}\}$ and at least two of each set $\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\}$, for $2\leq n\leq 26$, are transcendental, with analogous results for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$. We then show, via a theorem of Shidlovskii, that the $\eta^{\left(n\right)}$ are algebraically independent, and therefore transcendental, for all $n\geq 0$, implying that at least one of each pair, $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\}$ and $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\}$, and at least two of the triple $\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\}$, are transcendental for all $n\geq 1$. Further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ reveals that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often with positive asymptotic density. Finally, we provide parallel results for the sequences $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$ satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$.
@@ 1.3. Outline of Results
In Section 2, we begin our study of the arithmetic properties of the three generalized-constant sequences by investigating recurrence relations associated with the $\gamma^{\left(n\right)}$. These recurrences reveal that at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental. As will be seen, the latter result follows from verification, for $n\in\left\{2,3,\ldots,26\right\}$, of a certain reasonable conjecture (which, if true for all $n\in\mathbb{Z}_{\geq 2}$, necessitates that the $\gamma^{\left(n\right)}$ are transcendental infinitely often). Since $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$, all conclusions for the $\gamma^{\left(n\right)}$ apply equally well to the sequence $\Gamma^{\left(n\right)}\left(1\right)$.
In Section 3, we employ a theorem of Shidlovskii (1989) to demonstrate the algebraic independence, and therefore transcendence, of the $\eta^{\left(n\right)}$ for all $n\geq 0$. This implies that at least one element of each pair, $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e\right\}$ and $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}\right\}$, and at least two elements of the triple $\left\{\gamma^{\left(n\right)},\delta^{\left(n\right)}/e,\delta^{\left(n\right)}\right\}$, are transcendental for all $n\geq 1$. Moreover, further analysis of the $\gamma^{\left(n\right)}$ and $\eta^{\left(n\right)}$ shows that the values $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$. Finally, Section 4 addresses two additional sequences, $\widetilde{\delta}^{\left(n\right)}$ and $\widetilde{\eta}^{\left(n\right)}$, satisfying the “non-alternating analogue” equation $\gamma^{\left(n\right)}+\widetilde{\delta}^{\left(n\right)}/e=\widetilde{\eta}^{\left(n\right)}$, and provides results parallel to those for the original system.
@@ 2.2.1. Algebraic Structure of the $P_{n}$
Assume (for purposes of contradiction) that the polynomials $\left\{P_{2},P_{3},\ldots,P_{n}\right\}$ are algebraically dependent; that is, there exists a non-zero polynomial $Q_{n}\in\mathbb{Q}\left[t_{2},t_{3},\ldots,t_{n}\right]$ such that $Q_{n}\left(P_{2},P_{3},\ldots,P_{n}\right)=0$. [^8] If $Q_{n}$ does not actually involve the variable $t_{n}$, then $Q_{n}\in\mathbb{Q}\left[t_{2},t_{3},\ldots,t_{n-1}\right]$ and $Q_{n}\left(P_{2},P_{3},\ldots,P_{n}\right)=0$ already embodies an algebraic dependence among $\left\{P_{2},P_{3},\ldots,P_{n-1}\right\}$. In that case, we can replace $n$ by n1n-1 (and repeat this reduction if necessary) so that the assumed dependence relation involves the highest-index variable, $n^{*}$. Then, without loss of generality, we take $n=n^{*}$. Writing $Q_{n}$ as a polynomial in $t_{n}$ gives
@@ 2.3. Transcendence Results
The transformation of the original polynomials of (16) into the reduced form of (17) lays the foundation for part (ii) of Theorem 1 below. However, one additional condition on the $\widetilde{P}_{n,\theta}$ is required. Although we know there exists $\widetilde{P}_{n,\theta}\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)=0$ that can be expressed as a non-trivial polynomial equation in $\theta\in\left(\mathcal{G}_{n}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)$ of the form given by (19), it remains possible that $A_{d}^{\left(n,\theta\right)}=0$ for all $d\in\left\{1,2,\ldots,D_{n,\theta}\right\}$ at the actual realizations of the elements $\gamma^{\left(k\right)}\in\mathcal{G}_{n}^{\left(2n\right)}$ (to be denoted by $\boldsymbol{\gamma^{*}}\Longleftrightarrow\gamma^{\left(k\right)}=\textrm{E}_{X}\left[X_{\textrm{G}}^{k}\right]\>\forall\>k\in\mathbb{Z}_{\geq 1}$). Lemma 5 below (supported by the subsequent Conjecture 1) ensures this does not happen.
@@
The conjecture below, which we have verified computationally for all $n\in\left\{2,3,\ldots,26\right\}$, asserts that condition (21) of Lemma 5 is true. [^9] For each $n$, we executed a Python program (on an Apple M4 Max MacBook computer with 48 GB memory) to provide rigorous certification of the sufficient condition in (21). The program first evaluated $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$ symbolically for each $m\in\left\{2,3,\ldots,n\right\}$, and then confirmed $\left.\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)\right|_{\boldsymbol{\gamma^{*}}}\neq 0$ using interval arithmetic. The maximum value of $n$ tested (i.e., $26$) was simply the largest $n$ for which the program completed all checks within approximately 24 hours. Clearly, greater values of $n$ could be tested, but the program’s runtime increases approximately exponentially in $n$. (For example, $n=30$ would require a runtime of about 1 week.) In addition to its computational verification, the conjecture is realistic on theoretical grounds because it holds for any given $n$ unless $\boldsymbol{\gamma^{*}}$ is forced to lie on one of finitely many special hypersurfaces (i.e., the vanishing loci of the minors $\det\left(\mathbf{J}_{m,\theta}^{\left(n\right)}\right)$).
@@
(iv) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ (i.e., $\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $\Omega\left(\ln\left(N\right)/N\right)$ as $N\rightarrow\infty$.
@@
(iv) As previously observed, if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then part (ii) forces each set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ to contain at least one transcendental element $\gamma^{\left(k_{n}\right)},\>k_{n}\in\left\{n,n+1,\ldots,2n\right\}$. Taking $m=3^{i}$ for $i\in\mathbb{Z}_{\geq 1}$, we now consider the family of disjoint intervals $\left[3^{i},2\cdot 3^{i}\right]\ni k_{m}$ and note that we can choose $N\in\mathbb{Z}_{\geq 2}$ such that $N\geq 2\cdot 3^{i}=2m$ and the $k_{m}\in\left[3^{i},2\cdot 3^{i}\right]\subseteq\left\{1,2,\ldots,N\right\}$ are distinct. This implies that the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N$ is given by
$\dfrac{\#\left\{n\leq N:\gamma^{\left(n\right)}\textrm{ is transcendental}\right\}}{N}\geq\dfrac{\#\left\{k_{m}\leq N:\gamma^{\left(k_{m}\right)}\textrm{ is transcendental}\right\}}{N}$
$\geq\dfrac{1}{N}\left\lfloor\log_{3}\left(\dfrac{N}{2}\right)\right\rfloor\geq\dfrac{1}{N}\left(\dfrac{\ln\left(N\right)-\ln\left(2\right)}{\ln\left(3\right)}-1\right),$
with asymptotic lower bound $\Omega\left(\ln\left(N\right)/N\right)$ as $N\rightarrow\infty$. ∎
Since Conjecture 1 has been verified computationally for $n\in\left\{2,3,\ldots,26\right\}$, we know from part (ii) of the above theorem that at least two elements of each associated set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental. Furthermore, it is easy to see from the $\textrm{Gumbel}\left(0,1\right)$ moment-generating function, $\textrm{M}_{X}\left(t\right)=\Gamma\left(1-t\right),\>t<1$, that
@@
implying that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere. In particular: (a) at least two elements of each set $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(n\right)}\left(1\right),\Gamma^{\left(n+1\right)}\left(1\right),\ldots,\Gamma^{\left(2n\right)}\left(1\right)\right\}$ are transcendental for $n\in\left\{2,3,\ldots,26\right\}$; and (b) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\Gamma^{\left(n\right)}\left(1\right)$ must be transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]). [^10] It may be of interest to some readers to note that the asymptotic lower bound on the density of the set of posited transcendental $\Gamma^{\left(n\right)}\left(1\right)$ is smaller (in magnitude) than the asymptotic lower bound on the density of the set of irrational odd-$n$ $\zeta\left(n\right)$ given by Fischler, Sprang, and Zudilin [12]: $\Omega\left(\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{N/2}\right)=\Omega\left(N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1}\right),$ for arbitrarily small $\varepsilon>0$.
@@ 3. Algebraic Independence of the $\eta^{\left(n\right)}$
From (16), one can see that for any $m\in\left\{2,3,\ldots,n\right\}$, $P_{m}$ is linear in the highest-index constant, $\gamma^{\left(2m\right)}$, with a non-zero integer coefficient $\alpha_{m,0}$. This implies that each $\gamma^{\left(2m\right)}$ can be expressed as a polynomial with rational coefficients in strictly lower-index constants; that is, $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}\right)$. We now let
@@
For $m=2$, the relation $P_{2}=0$ expresses $\gamma^{\left(4\right)}$ as a polynomial with rational coefficients in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)}\in\mathcal{B}_{n}$, implying $\gamma^{\left(4\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$. Now consider $m\geq 3$, and assume that $\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$. Since (a) $P_{m}=0$ expresses $\gamma^{\left(2m\right)}$ as a polynomial with rational coefficients in $\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(2m-1\right)}$, (b) every odd-indexed element of the set $\left\{\gamma^{\left(3\right)},\gamma^{\left(5\right)},\ldots,\gamma^{\left(2m-1\right)}\right\}$ lies in $\mathcal{B}_{n}$, and (c) every even-indexed element of $\left\{\gamma^{\left(4\right)},\gamma^{\left(6\right)},\ldots,\gamma^{\left(2m-2\right)}\right\}$ lies in $\mathbb{Q}\left(\mathcal{B}_{n}\right)$, it follows that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ as well. This confirms that $\gamma^{\left(2m\right)}\in\mathbb{Q}\left(\mathcal{B}_{n}\right)$ for all $m\in\left\{2,3,\ldots,\lfloor n/2\rfloor\right\}$, and thus that $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}=\mathbb{Q}\left(\mathcal{G}_{n}\right)=\mathbb{Q}\left(\mathcal{B}_{n}\right)$.
@@
Thus, as $N\rightarrow\infty$, the density has asymptotic lower bound